# How Memory Management Impacts LLM Agents: An Empirical Study of Experience-Following Behavior

### v1

@@ Abstract
Memory is a critical component in large language model (LLM)-based agents, enabling them to store and retrieve past executions to improve task performance over time. In this paper, we conduct an empirical study on how memory management choices impact the LLM agents’ behavior, especially their long-term performance. Specifically, we focus on two fundamental memory operations that are widely used by many agent frameworks—addition, which incorporates new experiences into the memory base, and deletion, which selectively removes past experiences—to systematically study their impact on the agent behavior. Through our quantitative analysis, we find that LLM agents display an experience-following property: high similarity between a task input and the input in a retrieved memory record often results in highly similar agent outputs. Our analysis further reveals two significant challenges associated with this property: error propagation, where inaccuracies in past experiences compound and degrade future performance, and misaligned experience replay, where outdated or irrelevant experiences negatively influence current tasks. Through controlled experiments, we show that combining selective addition and deletion strategies can help mitigate these negative effects, yielding an average absolute performance gain of 10% compared to naive memory growth. Furthermore, we highlight how memory management choices affect agents’ behavior under challenging conditions such as task distribution shifts and constrained memory resources. Our findings offer insights into the behavioral dynamics of LLM agent memory systems and provide practical guidance for designing memory components that support robust, long-term agent performance. We also release our code to facilitate further study.  [^1] The dataset and experimental code are available in https://github.com/yuplin2333/agent_memory_manage.git
@@ 1 Introduction
Figure 1: Left: Illustration of the memory management workflow after each agent execution. Right: identified challenges along with experience following, including error propagation (Section 4.4): errors in previously stored executions propagate to similar tasks when being retrieved; and misaligned experience replay (Section 5.3): similar but inherently misaligned records weaken experience-following effectiveness.
Recently, advances in Large Language Models (LLMs) with strong instruction-following capabilities have paved the way for LLM agents wang2024survey ; zhang2024survey ; xi2023rise – an autonomous system powered by LLM that can interact with the external environment and perform tasks. Unlike stand-alone LLMs, LLM agents possess the ability to perceive the environment, engage in step-by-step reasoning and planning, evolve through memory, and make decisions to solve complex real-world tasks wu2023autogen ; schmidgall2025agent ; wang2023voyager ; shinn2024reflexion ; yao2023react .
Among the various components, memory plays a pivotal role in LLM agent execution wang2024survey . It retains past task queries (i.e. the agent input) and execution (i.e. the agent output), which can be retrieved as in-context demonstrations to guide similar future tasks. Through effective memory management, including adding, updating, and deleting past experiences, the performance of LLM agents can be improved over time zhang2024survey . However, despite the importance, the properties of the general memory module for LLM agents and their management remain largely under-explored. While some studies have investigated effective memory management for specific agent types yin2024explicit ; zeng2024structural ; zhao2024expel ; zhong2024memorybank ; wang2024agent , they fail to provide a broader understanding that could inform the development of better memory systems for a wide range of autonomous LLM agents. The major obstacle to an in-depth study of the memory module lies in the diversity of agents’ task objectives and functionalities, which leads to highly varied–and often inconsistent–memory designs  shi2024ehragent ; mao2023language ; wang2023voyager ; li2023tradinggpt ; park2023generative ; xiang2024guardagent . For example, EHRAgent shi2024ehragent stores generated code snippets for tasks related to electronic health records (EHRs), whereas AgentDriver mao2023language maintains a vector encoding past states and corresponding trajectory prediction by the agent.
To bridge this gap, we present an in-depth quantitative study to uncover the fundamental principles of memory management for general autonomous LLM agents. To draw broadly applicable insights across diverse memory designs, we focus on two essential memory operations–addition and deletion–which are widely adopted by existing agents, as illustrated in Figure 1. We quantitatively investigate how different memory addition and deletion designs shape agent behavior and identify an important phenomenon that we term the experience-following property: a high ‘input similarity’ between the current task query and the one from the retrieved record often yields a high ‘output similarity’ between their corresponding (output) executions.
@@
While the experience-following property supports effective reuse of successful experiences, we also uncovered two notable challenges associated with it. First, we observe the problem of error propagation: if a retrieved memory record contains low-quality or incorrect outputs, the agent is likely to replicate and even amplify these errors during the current task. If the resulting execution is then added back into memory, the error is likely to be further propagated to future tasks (Figure 1). Second, we recognize the issue of misaligned experience replay, which limits the benefits of experience following–certain memory records, when retrieved as demonstrations, consistently result in low output similarity, showing their inadequacy as demonstrations. Retaining these records increases the likelihood of suboptimal or incorrect executions (Figure 1).
Through controlled experiments on three agents across diverse tasks, we systematically analyze how selective strategies for the addition and deletion of memory entries can influence long-term agent behavior. Our findings show that to mitigate error propagation, a Selective Addition approach should be employed, where each task query and its associated output execution are assessed for quality insurance before being integrated into the memory. Furthermore, to counter misaligned experience replay and reduce memory size for more efficient storage, we propose a simple Combined Deletion strategy: redundant records will be deleted based on recent retrieval frequencies, while low-quality and outdated records will be deleted based on their historical task utilities as demonstrations.
@@
We further validate our observations under challenging conditions representing real-world scenarios, including (1) Task Distribution Shift: the task distribution changes substantially over time, requiring the agent to adapt to shifting patterns and contexts. (2) Memory Resources Constraint: the memory capacity is severely limited, requiring the agent to retain only the most valuable and helpful experience. The results underscore the importance of thoughtful memory management for maintaining stable and effective long-term agent performance. Our contributions are summarized below:
@@
  • We conduct systematic and quantitative analysis of the impact of the two most essential memory operations—addition and deletion—of the memory records on the long-term performance of LLM agents. Our findings provide key principles for robust memory management design.
@@
  • We propose and advocate a general framework incorporating a selective addition and a combined deletion policy. Through the evaluation of three agents for three different tasks, respectively, we show the effectiveness of this framework in mitigating both error propagation and misaligned experience replay, with an average 10% absolute performance gain over the baseline policies.
  • We further validate the necessity of these policies under two demanding real-world conditions: task distribution shift and memory resource constraints. Our findings show that the combined policy helps agents adapt to evolving tasks and operate effectively with limited memory.
@@ 2.1 Memory Module of LLM Agents
LLM agents often include short-term memory and long-term memory zhang2024survey . Short-term memory usually refers to inside-task working memory sumers2023cognitive , while long-term memory sumers2023cognitive can be divided into three types: semantic memory, procedural memory, and episodic memory. Semantic memory Kumar2020SemanticMA contains the agent’s world knowledge and understanding of the environment; procedural memory proceduralmemory involves rules or procedures, which may reside implicitly in the LLM’s weights or be explicitly defined as guidelines for the agent; and episodic memory Nuxoll2007ExtendingCA records task-specific experiences. In this paper, we focus on episodic memory, which is commonly used by LLM agents with memory.
@@ 2.2 Memory Management and Limitations
LLM agents often employ an episodic memory module to store past experiences for future retrieval, which is crucial for effective planning and task execution shi2024ehragent ; mao2023language ; shinn2024reflexion ; hong2023metagpt ; wang2023voyager . The memory module typically includes operations such as memory reading and memory management zhang2024survey . Specifically, given a new task query $q$ and a memory base currently containing $N$ query-execution pairs $\mathcal{D}=\{(q_{1},e_{1}),\ldots,(q_{N},e_{N})\}$, the agent’s execution cycle involves the following memory operations:
@@
The agent retrieves a subset $\xi_{K}\subset\mathcal{D}$ consisting of the $K$ query-execution pairs most relevant to the query $q$. The relevance is often measured by the input similarity between the task query $q$ and the query from the retrieved past experience. For example, the relevance can be computed as the cosine similarity between the feature representations for both queries from a text encoder kusupati2022matryoshka . The retrieved pairs $\xi_{K}$ are then used as in-context learning demonstrations to guide the LLM in generating an execution trajectory $e$ for the task query $q$.
@@
Memory management typically involves the addition and deletion of memory records. Specifically, when obtaining the query-execution trajectory pair $(q,e)$, addition decides whether this pair should be added to memory, while deletion zhong2024memorybank ; liu2023think is often triggered to remove outdated or redundant pairs from memory.
Although various strategies have been proposed for memory management, such as structural transformation wang2023voyager ; anokhin2024arigraph ; zeng2024structural , merging zhong2024memorybank ; liu2023think ; hu2024hiagent , summarization wang2024agent ; zhong2024memorybank , and reflection shinn2024reflexion ; zhao2024expel , these approaches are often designed for specific agent types (e.g., chatbot agents) and do not provide a unified set of principles. Parallel works also conduct quantitative evaluations of different memory management methods, including a structural transformation and retrieval method zeng2024structural and an expectation maximization approach for optimizing the memory bank yin2024explicit . However, their findings remain tailored to certain agent types and are difficult to extend to more complex LLM agents, such as code agents covered by our study yin2024explicit .
In contrast, our work centers on the basic operations of memory addition and deletion, aiming to uncover how these basic memory operations influence the behavior and long-term performance of generic LLM agents. Our findings will shed light on potential challenges posed by these operations and offer insights to guide future memory system design.
3 Evaluation Setup
In this paper, we investigate memory management in generic LLM agent settings by considering three representative agents designed for different tasks: EHRAgent shi2024ehragent , AgentDriver mao2023language , and CIC-IoT Agent ciciot . In addition to their distinct tasks, these agents also vary significantly in input-output formats and memory retrieval mechanisms, which enhances the generalizability of our findings. See Table 3 in Appendix A.1 for a high-level summary of these agents.
EHRAgent shi2024ehragent : EHRAgent is a code-generation agent that allows clinicians to interact with electronic health records (EHRs) using natural language queries. We adopt the MIMIC-III dataset 2016NatSD…360035J for evaluation. By filtering out duplicated data and those lacking answers, we get 2392 tasks in total. Following the original study, we retrieve four past experiences during the execution of each test query based on a general text embedding model and maximum cosine similarity. The initial memory bank contains 100 records. The accuracy (ACC) is reported as task performance.
AgentDriver mao2023language AgentDriver is an LLM-based autonomous driving agent that integrates common sense and experience into the memory record. Following the original setup, we use the nuScenes dataset caesar2020nuscenes and randomly sample 2000 test cases from its test set. We slightly simplify the original two-step retrieval of AgentDriver to a single-step retrieval using top-1 vector similarity, i.e., the retrieval results in a single demonstration (see Appendix A.2 for details). Task performance is evaluated using the success rate (SR), which measures the proportion of executions that predict a trajectory with an average L2 error below 2.5. Here, the L2 error is measured in UniAD hu2023planning over 3-second, which quantifies the distance between the predicted and ground-truth trajectories. For the initial memory, we randomly sample 180 experiences from the nuScenes training set.
CIC-IoT Agent CIC-IoT Agent is a network intrusion detection agent designed to predict attack types based on IoT packet features. Specifically, we construct prompts by formatting these features according to a predefined template and input them into the Agent, which then predicts the corresponding attack type. We evaluate CIC-IoT Agent using the CIC-IoT benchmark dataset ciciot . Given the large scale of the dataset, we randomly sample 1200 test cases from the test set. For each test query, we retrieve three past experiences during execution using a feature-based approach, as detailed in Appendix A.3. To initialize the memory bank, we use GPT-4o to generate 100 records from a disjoint training set. Task performance is measured by accuracy (ACC)
For all three agents, we use GPT-4o-mini as the backbone language model for most of the experiments. More details can be found in Appendix A.
@@ 4 Addition of Memory
Memory addition refers to the process of deciding whether a finished task execution should be stored as a new record in the memory bank. We investigate three memory addition strategies—an add-all baseline, selective addition based on coarse automatic evaluation, and selective addition based on strict human evaluation—along with a fixed-memory baseline without memory addition. Through our empirical evaluation of these strategies, we have three central findings: (1) Selection is the key to effective memory addition: Selective addition based on strict human evaluation consistently surpasses the fixed-memory baseline and the add-all strategy without a memory filter (Section 4.2). (2) Agent execution exhibits an experience-following property: When the retrieved task query closely resembles the current task, the input similarity often strongly correlates with the similarity of their output executions (Section 4.3). (3) Error propagation: Through an in-depth investigation of the experience-following property, we identify an error propagation challenge for memory management and demonstrate how selective addition helps mitigate this issue (Section 4.4).
4.1 Setup for Memory Addition Experiments
@@
We begin by describing several memory addition strategies. For a query-execution pair $(q,e)$ and an evaluator $\pi$, the addition decision is given by $\pi(q,e)$: if $\pi(q,e)=1$, the experience is stored; if $\pi(q,e)=0$, it is discarded. Our aim is to analyze how different addition strategies influence the agent’s behavior and performance in the long run. Starting from an identical initial memory, we examine the following four memory addition strategies:
1) Fixed-memory baseline:
@@
To establish a baseline and demonstrate the need for memory addition, we first consider a fixed memory bank. In this setting, we use a subset of the training data with ground truth agent execution as the fixed memory bank mao2023language ; li2023tradinggpt ; zhao2024expel , and the agent relies only on this fixed memory without adding new entries, i.e., $\pi_{\rm fixed}(q,e)=0$.
2) Add-all approach:
@@
A straightforward baseline is to store every encountered task and its execution: $\pi_{\rm all}(q,e)=1$.
3) Selective addition based on automatic evaluation (coarse):
Coarse selective addition employs an LLM to decide whether an execution should be added to the memory, i.e., $\pi_{\rm automatic}(q,e)=\text{LLM}(q,e)$. This method is broadly applicable and relatively cost-effective. Detailed designs of these automatic evaluators are provided in Appendix A.4.
4) Selective addition based on human evaluation (strict):
@@
In this stricter approach, a human (oracle) serves as the evaluator, determining whether the execution should be stored in the memory: $\pi_{\rm human}(q,e)=\text{Human}(q,e)$. Although human input could be gathered after each execution for practical agents, it is not feasible for our current evaluation. Therefore, we simulate this process by comparing the generated output with the ground truth.
Details on each agent’s specific designs for memory addition can be found in Appendix A.2.
4.2 Selection is Necessary for Memory Addition
@@
Table 1: Performance of the three memory addition strategies, add-all: selective addition based on automatic evaluation (coarse), and selective addition based on human evaluation (strict), compared with the fixed-memory baseline (shaded) on EHRAgents, AgentDriver, and CIC-IoT. The results highlight the necessity of selection for effective memory addition.
Strategy EHRAgents AgentDriver CIC-IoT Agent
ACC $\uparrow$ Mem Size $\downarrow$ SR. $\uparrow$ Mem Size $\downarrow$ ACC. $\uparrow$ Mem Size $\downarrow$
Fixed 16.89 100 40.53 180 60.25 100
Add all 13.04 2411 32.48 2125 60.00 1300
Coarse 31.35 1881 37.03 1161 60.67 1083
Strict 38.86 1012 50.94 1178 63.33 860
We examine the performance of selective addition, add-all, and fixed-memory strategies. First, we compare their overall performance over the long run, and second, we analyze the trends in the performance and the underlying reasons.
Selective addition based on strict criteria outperforms all other strategies, while add-all is inferior to the fixed-memory baseline.
Table 1 summarizes the performance of the three memory addition strategies on the three agents, respectively. When comparing add-all with the fixed-memory baseline, we observe that simply storing every experience leads to significantly worse outcomes, particularly for EHRAgents and AgentDriver. In contrast, both coarse and strict selective addition maintain reasonable memory sizes while achieving higher performance in most cases compared with add-all.
Selective addition improves the long-term performance of agents through self-evolution, while add-all causes self-degradation.
Figure 2: Performance trend over AgentDriver. Strict addition consistently improves the performance over time.
Although the same initial memory is used by all methods in our evaluation, the long-term performances of these methods diverge considerably. Figure 2 and Figure 8 in Appendix B illustrate the accuracy trend for these three agents using each addition strategy. The curve for the add-all strategy either remains flat or declines over time, suggesting a probable accumulation of flawed entries that can degrade the agent’s performance. The curve for the fixed-memory baseline stays above the add-all approach but exhibits little improvement for EHRAgent and AgentDriver, as no examples could be further incorporated into the memory. In contrast, the performance for strict selective addition continues to improve over extended runs. This trend is also verified using other LLM backbones, as shown in Appendix C.1 These results underscore the advantages of selectively integrating new experiences.

### v2

@@ Abstract
Memory is a critical component in large language model (LLM)-based agents, enabling them to store and retrieve past executions to improve task performance over time. In this paper, we conduct an empirical study on how memory management choices impact the LLM agents’ behavior, especially their long-term performance. Specifically, we focus on two fundamental memory management operations that are widely used by many agent frameworks—memory addition and deletion—to systematically study their impact on the agent behavior. Through our quantitative analysis, we find that LLM agents display an experience-following property: high similarity between a task input and the input in a retrieved memory record often results in highly similar agent outputs. Our analysis further reveals two significant challenges associated with this property: error propagation, where inaccuracies in past experiences compound and degrade future performance, and misaligned experience replay, where some seemingly correct executions can provide limited or even misleading value as experiences. Through controlled experiments, we demonstrate the importance of regulating experience quality within the memory bank and show that future task evaluations can serve as free quality labels for stored memory. Our findings offer insights into the behavioral dynamics of LLM agent memory systems and provide practical guidance for designing memory components that support robust, long-term agent performance. We also release our code to facilitate further study.  [^1] The dataset and experimental code are available in https://github.com/yuplin2333/agent_memory_manage.git
How Memory Management Impacts LLM Agents:
An Empirical Study of Experience-Following Behavior
Zidi Xiong1thanks: Equal contributions. Correspondence to: Zidi Xiong zidixiong@g.harvard.edu and Zhen Xiangzhen.xiang.lance@gmail.com, Yuping Lin3∗, Wenya Xie4∗, Pengfei He3, Zirui Liu4, Jiliang Tang3, Himabindu Lakkaraju1, Zhen Xiang2 1Harvard University  2University of Georgia 3Michigan State University 4University of Minnesota-Twin Cities
@@ 1 Introduction
Figure 1: Illustration of the memory management workflow after each agent execution.
To enable effective solving of complex tasks and self-evolving over time, large language model (LLM) agents often equip an episodic memory module Wang et al. (2024a). This dynamic mechanism retains past task queries (i.e. the agent input) and execution (i.e. the agent output), which can be retrieved as demonstrations to guide similar future tasks. Through effective memory management, including adding, updating, and deleting past experiences, the performance of LLM agents can be improved over time Zhang et al. (2024).
In practice, memory designs vary widely across agents due to their diverse objectives and functionalities Shi et al. (2024); Mao et al. (2024); Wang et al. (2023); Li et al. (2023); Park et al. (2023); Xiang et al. (2024). To address this heterogeneity, recent studies have proposed a range of memory management strategies Yin et al. (2024); Zeng et al. (2024); Zhao et al. (2024); Zhong et al. (2024); Xu et al. (2025). However, these approaches are often tailored to specific tasks and offer limited understanding of the underlying principles that govern memory behavior across different agentic systems.
In this work, we investigate a fundamental question in memory management: How do the evolving dynamics of the memory bank, driven by continuous memory operations, influence long-term agent execution? In particular, we focus on the core mechanisms underlying these dynamics—memory addition and memory deletion—as they constitute the foundation of memory management (see Figure 1). Moreover, unlike prior studies on in-context learning with static, external knowledge bases (Luo et al., 2024), our setting with these two operations captures the distinctive nature of agentic memory systems: a dynamic retrieval pool (memory bank) that evolves over time and contains inherently noisy outputs (trajectories) that are often generated by the agent itself.
@@
Through extensive experiments, we identify an important phenomenon that we term the experience-following property: a high ‘input similarity’ between the current task query and the one from the retrieved record often yields a high ‘output similarity’ between their corresponding (output) executions. While this property enables effective reuse of successful experiences, we uncover two significant challenges arising from the dynamic and noisy nature of agentic memory banks. First, we observe the problem of error propagation: if a retrieved memory record contains noisy or incorrect outputs, the agent is likely to replicate and even amplify these errors during the current task. If the resulting execution is then added back into memory, the error is likely to be further propagated to future tasks. Second, we recognize the issue of misaligned experience replay, which limits the benefits of experience following—certain memory records, when retrieved as demonstrations, consistently lead to poor execution due to their misalignment with the current task, indicating their inadequacy as effective demonstrations. Retaining these records increases the likelihood of suboptimal or incorrect executions.
To systematically analyze these phenomena, we conduct controlled experiments across four distinct agents operating on diverse tasks. Crucially, we reveal the critical yet underexplored role of trajectory evaluators in both memory addition and deletion processes, providing novel insights related to effective memory bank optimization that have been significantly overlooked in existing literature.
@@
We further examine challenging conditions representing real-world scenarios, including (1) Task Distribution Shift: the task distribution changes substantially over time, requiring the agent to adapt to shifting patterns and contexts. (2) Memory Resources Constraint: the memory capacity is severely limited, requiring the agent to retain only the most valuable and helpful experience. The results underscore the importance of thoughtful memory management for maintaining stable and effective long-term agent performance.
Our contributions are summarized below:
@@
  • We conduct systematic and quantitative analysis of the agent memory dynamics through two most essential memory operations—addition and deletion—of the memory records on the long-term performance of LLM agents. Our findings provide key principles for robust memory management design.
@@
  • We demonstrate how incorporating environmental feedback from trajectory evaluators can effectively improve memory management by mitigating both error propagation and misaligned experience replay.
  • We study how memory evolution affects long-term agent performance under task distribution shifts and memory constraints, finding that well-chosen evaluators enable agents to adapt to these scenarios and sustain performance with simple addition and deletion.
@@ 2.1 Memory Module of LLM Agents
LLM agents often include short-term memory and long-term memory Zhang et al. (2024). Short-term memory usually refers to inside-task working memory Sumers et al. (2023), while long-term memory Sumers et al. (2023) can be divided into three types: semantic memory, procedural memory, and episodic memory. Semantic memory Kumar (2020) contains the agent’s world knowledge and understanding of the environment; procedural memory Beaunieux et al. (2006) involves rules or procedures, which may reside implicitly in the LLM’s weights or be explicitly defined as guidelines for the agent; and episodic memory Nuxoll and Laird (2007); Lampinen et al. (2025) records task-specific experiences. In this paper, we specifically focus on episodic memory, which is commonly used by LLM agents with memory and has recently been shown to mitigate the lack of latent learning capability in LLMs (Lampinen et al., 2025).
@@ 2.2 Memory Management and Limitations
LLM agents often employ an episodic memory module to store past experiences for future retrieval, which is crucial for effective planning and task execution Shi et al. (2024); Mao et al. (2024); Xu et al. (2025); Zhou et al. (2025); Zhong et al. (2024); Yin et al. (2024). The memory module typically includes operations such as memory reading and memory management Zhang et al. (2024). Specifically, given a new task query $q$ and a memory base currently containing $N$ query-execution pairs $\mathcal{D}=\{(q_{1},e_{1}),\ldots,(q_{N},e_{N})\}$, the agent’s execution cycle involves the following memory operations:
@@
The agent retrieves a subset $\xi_{K}\subset\mathcal{D}$ consisting of the $K$ query-execution pairs most relevant to the query $q$. The relevance is often measured by the input similarity between the task query $q$ and the query from the retrieved past experience. For example, the relevance can be computed as the cosine similarity between the feature representations for both queries from a text encoder Kusupati et al. (2022). The retrieved pairs $\xi_{K}$ are then used as in-context learning demonstrations to guide the LLM in generating an execution trajectory $e$ for the task query $q$.
@@
Memory management typically involves the addition and deletion of memory records. Specifically, when obtaining the query-execution trajectory pair $(q,e)$, addition decides whether this pair should be added to memory, while deletion Zhong et al. (2024); Wang et al. (2024b) is often triggered to remove outdated or redundant pairs from memory.
Although various strategies have been proposed for memory management, such as structural transformation Anokhin et al. (2024); Zeng et al. (2024); Xu et al. (2025), merging Zhong et al. (2024); Liu et al. (2023); Hu et al. (2024), summarization Wang et al. (2025); Zhong et al. (2024); Pan et al. (2025), and reflection Shinn et al. (2024); Zhao et al. (2024); Liang et al. (2024), these approaches are often designed for specific agent types (e.g., chatbot agents) and lack a unified design that generalizes across different agentic systems.
Recent works also study the optimization of the memory bank, such as using an expectation–maximization approach for optimization Yin et al. (2024) or formulating memory management as a Markov Decision Process (Zhou et al., 2025). However, these works do not provide an understanding of how memory bank dynamics affect long-term performance under the inherently noisy nature of basic addition and deletion operations. In contrast, our findings highlight potential challenges posed by these operations and offer insights to guide the design of future memory systems.
@@ 3 Addition of Memory
Memory addition refers to the process of deciding whether a finished task execution should be stored as a new record in the memory bank. Here, we consider the most common approach that adds memory based on the feedback of the trajectory evaluator (Wang et al., 2025; Shi et al., 2024). We investigate three types of memory addition with different trajectory evaluators—an add-all baseline, selective addition based on coarse automatic evaluation, and selective addition based on strict human evaluation—along with a fixed-memory baseline without memory addition.
3.1 Setup
Setup for Agents
We investigate memory management in generic LLM agent settings by considering an illustrative and controllable synthetic agent and three representative agents designed for different tasks: For the synthetic agent, we design a task in which the agent predicts the output of a simple mapping – a linear function in the ground truth – by performing regression approximation over the provided demonstrations, which we call RegAgent. Formally, the agent receives an input vector $x$ together with several past guesses over nearby inputs under the linear transformation $w$ to infer an output for $w^{\top}x$. This setup allows us to investigate an ideal setting where the model fully relies on the provided demonstrations to generate the output, while enabling control over the noise level of the stored memory and direct measurement of the resulting error.
For real agent, we include EHRAgent Shi et al. (2024), AgentDriver Mao et al. (2024), and CIC-IoT Agent Neto et al. (2023). In addition to their distinct tasks, these agents also vary significantly in input-output formats and memory retrieval mechanisms, which enhances the generalizability of our findings. For all four agents, we use GPT-4o-mini as the backbone language model for most of the experiments. See Appendix A.1 for details of these agents.
Setup for Memory Addition
@@ 3.1 Setup
We begin by describing several memory addition using different trajectory evaluator. For a query-execution pair $(q,e)$ and an evaluator $\pi$, the addition decision is given by $\pi(q,e)$: if $\pi(q,e)=1$, the experience is stored; if $\pi(q,e)=0$, it is discarded. Our aim is to analyze how the noise introduced by additions from different evaluators influences the agent’s behavior and long-term performance. Starting from an identical initial memory, we examine the following four categories of memory additions:
@@
1) Fixed-memory baseline: In this baseline setting, we use a subset of the training data with correct agent execution as the fixed memory bank Mao et al. (2024); Li et al. (2023); Zhao et al. (2024); Zheng et al. (2024), and the agent relies only on this fixed memory without adding new entries, i.e., $\pi_{\rm fixed}(q,e)=0$.
@@
2) Add-all approach: A straightforward addition is to store every encountered task and its execution: $\pi_{\rm all}(q,e)=1$.
3) Selective addition based on automatic evaluation (coarse): We have $\pi_{\rm automatic}(q,e)=\text{Automatic}(q,e)$ and evaluate three automatic evaluators (Wang et al., 2025; Yin et al., 2024; Liang et al., 2024), denoted as Coarse 1 (C1), Coarse 2 (C2), and Coarse 3 (C3). For RegAgent, we define the evaluator based on the absolute error between the predicted and ground-truth outputs, where a smaller error threshold indicates a more accurate evaluator. Specifically, we set the thresholds to 1.6, 1.4, and 1.2 for C1, C2, and C3, respectively. For the other three agents, C1 and C2 correspond to GPT-4o-mini and GPT-4.1-mini used as evaluators, respectively, while C3 employs GPT-4.1-mini fine-tuned on 300 correct judge data collected from executions on a separate training set. Detailed designs of these automatic evaluators are provided in Appendix A.4.
@@
4) Selective addition based on human evaluation (strict):In this stricter approach, a human (oracle) serves as the evaluator (Shi et al., 2024), determining whether the execution should be stored in the memory: $\pi_{\rm human}(q,e)=\text{Human}(q,e)$. Although human input could be gathered after each execution for practical agents, it is not feasible for our current evaluation. Therefore, we simulate this process by comparing the generated output with the ground truth. More details on each agent’s specific designs for memory addition can be found in Appendix A.
3.2 Execution quality and memory size jointly determine long-term agent performance.
@@
Table 1: Performance of the three memory addition strategies: add-all, coarse evaluation with three different automatic evaluators, and selective addition based on human evaluation (strict), compared with the fixed-memory baseline (shaded). The results highlight the necessity of selection for effective memory addition.
Judge RegAgent EHRAgents AgentDriver CIC-IoT Agent SR. $\uparrow$ Mem Size $\downarrow$ ACC $\uparrow$ Mem Size $\downarrow$ SR. $\uparrow$ Mem Size $\downarrow$ ACC. $\uparrow$ Mem Size $\downarrow$ Fixed 67.53 100 16.75 100 40.11 180 71.50 50 Add all 55.48 4100 13.05 2411 32.32 2125 59.90 1050 Coarse C1 63.18 3511 26.19 1447 36.92 1161 74.00 1030 C2 65.78 3347 32.21 1467 40.01 1119 68.80 936 C3 67.35 3139 34.66 1094 47.37 1285 79.50 952 Strict Strict 70.95 2938 38.50 1012 51.00 1178 85.40 904
Figure 2: Performance trend for EHRAgent and AgentDriver. 4o-mini, 4.1-mini, and 4.1-mini FT denote different coarse evaluators from the GPT series. Both the strict evaluator and some coarse evaluators exhibit consistent self-improvement over time.
Figure 3: Left: Output similarity versus input similarity for RegAgent over different evaluators. Right: Output similarity versus input similarity for AgentDriver over different evaluators.
Figure 4: Comparison of running performance between using the agent output as demonstrations and the error-free (EF) variant using ground-truth. Coarse here uses C1 evaluator.
Table 1 summarizes the performance of memory addition strategies guided by five evaluators across four agents. Although all methods start with the same initial memory, their long-term performances diverge notably. The fixed-memory baseline maintains strong results with RegAgent, AgentDriver, and CIC-IoT Agent—sometimes even surpassing runs guided by coarse evaluators—indicating that noisy or low-quality additions can harm memory utility. In contrast, using a strict evaluator that selectively expands memory with high-quality records consistently yields superior performance, showing that both memory quality and capacity are key to effective long-term learning.
Furthermore, performance relies on the judge’s capability for coarse evaluators. For the three model judged agents, fine-tuning the evaluator on only 300 trajectories (C3) already achieves strong long-term improvements, outperforming other coarse evaluators and unfiltered additions.
As shown in Figure 2 and Figure 8 in Appendix B.1, the add-all and some coarse addition approach exhibits flat or declining success rates, suggesting accumulation of flawed records. The fixed-memory baseline performs better but remains static due to its frozen memory. In contrast, strict selective addition and C3 coarse evaluator continue to improve over time, a trend consistent across different LLM backbones (Appendix B.2) with strict addition.
These results suggest that careful construction of the evaluator is particularly important, as directly applying a vanilla LLM as a trajectory evaluator (Wang et al., 2025; Pan et al., 2024) may lead to a more severe negative impact than manually crafting a small but high-quality dataset.